the intensity I of light varies inversely as the square of the distance D from the source if the intensity of Illumination on a screen 5 ft from a light is 3 foot_candles find the intensity on a screen 15 ft from the light.
A. 1(1/3) Foot-candles
B.1/3 Foot-Candles
c.1/4 Foot- Candles
D. 2 Foot - Candles

Found 2 solutions by rothauserc, greenestamps:
Answer by rothauserc(4718)

I = k/D^2, where I is intensity and D is distance from source
:
Note k is the constant of proportionality
:
3 = k/5^2
:
k = 3 * 25 = 75
:
I = 75/15^2 = 75/225 = 1/3 foot_candle
:
answer is B.
:

Answer by greenestamps(13200)



Solving the problem as the other tutor did, using the formal definition of inverse proportion to find the constant of proportionality, is of course a valid method.

And you should be able to use that method, as it may be required in a more complicated problem.

But for a simple problem like this, I personally find it far easier simply to say that since the distance is multiplied by 3, the intensity is multiplied by 1/(3^2) = 1/9.

Then 1/9 of 3 foot-candles is 1/3 foot-candle, answer B.